# AAD-LLM: Neural Attention-Driven Auditory Scene Understanding

### v2

@@ Abstract
, Vishal Choudhari1, 3, Stephan Bickel4, 5,
Ashesh Mehta4, 5, Guy M McKhann2, Daniel Friedman6,Adeen Flinker6, Nima Mesgarani1, 3
@@
![[Uncaptioned image]](extracted/6281829/figures/ear.png)
![[Uncaptioned image]](extracted/6281829/figures/brain.png)
@@ 3.1 Intention Decoding
AAD-LLM introduces a speaker-based approach for decoding auditory attention from intracranial EEG signals. Instead of reconstructing speech features, we classify which speaker the listener is attending to using a discrete token representation. The process consists of two steps. First, we perform speaker clustering by applying K-means clustering to x-vectors Snyder et al. (2018) extracted from a large corpus of thousands of speakers, ensuring no overlap with test speakers. The x-vectors, commonly used in speaker verification, are 512-dimensional embeddings, which remain frozen during training. The number of clusters K is set to 8. Next, we perform speaker prediction from neural signals. A bidirectional LSTM maps the neural signal to a predicted cluster index . The ground-truth label is determined by finding the closest centroid to the attended speaker’s x-vector. Further details about the model is included in Appendix B.
@@ 4.3.2 Psychophysics Evaluation
Kruskal-Wallis H tests revealed significant differences in ratings between the groups for both Summarization and Free Q&A tasks (p-values < 0.001)11footnotetext: Asterisks indicate significance levels: * p0.05, ** p0.01, *** p0.001, and ‘ns’ denotes non-significant results (p0.05).. These tests were then followed up with post-hoc pairwise Bonferroni-corrected Mann-Whitney U tests. The results of these tests show that AAD-LLM’s responses were rated significantly higher than those from the mixture and random speaker baselines across both the tasks. While mean performance increased from brain-decoded to oracle attention with ground-truth speaker labels, this improvement was not statistically significant, indicating that the neural decoding is close to its best capacity. Furthermore, both AAD-LLM models (brain-decoded and oracle attention) approach the “oracle speaker” upper bound, suggesting that AAD-LLM’s responses closely mirror human perception. Please see Appendix F for more details on the evaluation methods, Tables 11 and 12 for p-values.
@@ D.3 Evaluation on Speech-only Dataset
D.4 Ablations
To better understand the contributions of different components and design choices in AAD-LLM, we conduct an ablation study with results presented in Table 9. We investigate the impact of the LoRA rank, the necessity of core components, and the type and clustering of features used for attention decoding.

### v3

@@ Abstract
, Vishal Choudhari1, 3,
Stephan Bickel4, 5,Ashesh Mehta4, 5, Guy M McKhann2,
Daniel Friedman6, Adeen Flinker6, Nima Mesgarani1, 3
@@
![[Uncaptioned image]](extracted/6530880/figures/ear.png)
![[Uncaptioned image]](extracted/6530880/figures/brain.png)
@@ 3.1 Intention Decoding
AAD-LLM introduces a speaker-based approach for decoding auditory attention from intracranial EEG signals. Instead of reconstructing speech features, we classify which speaker the listener is attending to using a discrete token representation. The process consists of two steps. First, we perform speaker clustering by applying K-means clustering to x-vectors Snyder et al. (2018) extracted from a large corpus of thousands of speakers, ensuring no overlap with test speakers. The x-vectors, commonly used in speaker verification, are 512-dimensional embeddings, which remain frozen during training. The number of clusters K is set to 8. Next, we perform speaker prediction from neural signals. A bidirectional LSTM maps the -channel -length neural signal to a predicted cluster index . The ground-truth label is determined by finding the closest centroid to the attended speaker’s x-vector. Further details about the model is included in Appendix B.
@@ 4.3.2 Psychophysics Evaluation
Kruskal-Wallis H tests revealed significant differences in ratings between the groups for both Summarization and Free Q&A tasks (p-values < 0.001)11footnotetext: Asterisks indicate significance levels: * p0.05, ** p0.01, *** p0.001, and ‘ns’ denotes non-significant results (p0.05).. These tests were then followed up with post-hoc pairwise Bonferroni-corrected Mann-Whitney U tests. The results of these tests show that AAD-LLM’s responses were rated significantly higher than those from the mixture and random speaker baselines across both the tasks. While mean performance increased from brain-decoded to oracle attention with ground-truth speaker labels, this improvement was not statistically significant, indicating that the neural decoding is close to its best capacity. Furthermore, both AAD-LLM models (brain-decoded and oracle attention) approach the “oracle speaker” upper bound, suggesting that AAD-LLM’s responses closely mirror human perception. Please see Appendix F for more details on the evaluation methods, Tables 12 and 13 for p-values.
@@ D.3 Evaluation on Speech-only Dataset
D.4 Impact of Background Noise Types
The clinical test set contains speech mixtures under two distinct background noises: babble and pedestrian. The performance with respect to the noise type is shown in Table 10. Transcription error (WER) and summarization quality (ROUGE-L & METEOR) are consistently better under babble noise compared to pedestrian noise. These results suggest that pedestrian noise, characterized by unpredictable acoustic dynamics, presents greater difficulty for the AAD-LLM, slightly degrading the recognition of attended speech and subsequent response generation.
NoiseType |
Babble | 87.1 | 86.5 | 10.7 | 29.0 | 59.5 / 56.5 | 44.8 / 43.6 | 65.5 / 68.4 | 63.1 / 63.5 | ||||||||||||||||||||||||
Pedestrian | 91.1 | 91.1 | 17.6 | 36.7 | 57.5 / 57.2 | 41.2 / 41.8 | 61.1 / 61.4 | 53.6 / 55.4 |
D.5 Impact of Window Size on Correct Speaker Selection
We also examined how varying the window size affects speaker selection accuracy, particularly in dynamic attention scenarios, with our speaker decoding method. The results, detailed in Figure LABEL:fig:window_size, demonstrate a clear trade-off: shorter windows offer rapid responsiveness but lower accuracy, whereas longer windows significantly improve decoding reliability at the expense of increased latency. These insights suggest the importance of adaptive strategies for window sizing tailored to specific real-time application demands.
D.6 Ablations
To better understand the contributions of different components and design choices in AAD-LLM, we conduct an ablation study with results presented in Table 9. We investigate the impact of the LoRA rank, the necessity of core components, and the type and clustering of features used for attention decoding.å